OpenStudy (theedwardsfamily):

What's 60+5 x 18-13?

OpenStudy (nothingwasthesame):

1157

OpenStudy (theedwardsfamily):

my answer was 137.

OpenStudy (theedwardsfamily):

so ur right! thnx

OpenStudy (doc.brown):

\[60+5\times 18+13\] \[60+(5\times 18)+13\]

OpenStudy (doc.brown):

\[60+90+13=163\]

OpenStudy (nothingwasthesame):

yea that would be 163 if its pemdas

OpenStudy (doc.brown):

It's not pemdas, it's math.

OpenStudy (anonymous):

why would you change the - to + 13

OpenStudy (nothingwasthesame):

no the order of pemdas i knows it math do you know what pemdas is ?

OpenStudy (doc.brown):

@Lcb2516 because math is easy, but LaTeX is difficult

OpenStudy (theedwardsfamily):

@doc.brown the way to work it out u use PEMDAS Please Excuse My Dear Aunt Sally

OpenStudy (nothingwasthesame):

yes you use pemdas

OpenStudy (nothingwasthesame):

i did not but thats what you use

OpenStudy (doc.brown):

\[60+5\times 18\color{red}{-}13=137\]

OpenStudy (theedwardsfamily):

double check ur wrk with a calcultor

OpenStudy (nothingwasthesame):

its 137

OpenStudy (nothingwasthesame):

you were right the first time

OpenStudy (doc.brown):

I did use a calculator, I just got the sign wrong. My apologies.

OpenStudy (nothingwasthesame):

@doc.brown good job though

OpenStudy (theedwardsfamily):

its ok

OpenStudy (anonymous):

:)